OpenStudy (anonymous):

write the equation 5x+y=4

OpenStudy (anonymous):

in slope intercept form

OpenStudy (uri):

Formula...? y = mx + b

OpenStudy (uri):

First 5x+y=4 y =-5x +4

OpenStudy (mathstudent55):

Solve the equation for y. Slope-intercept form is y = mx + b